# The calculus of Duistermaat’s triple index

### v2

@@ 4.2. The Maslov index
Suppose $\mathcal{L}\colon[0,1]\to\Lambda(n)$ is a differentiable path of Lagrangian planes, given by a frame $\big(X(t),Y(t)\big)$. For $t_{0}\in[0,1]$ and $v\in\mathcal{L}(t_{0})$ we define the [^5] The equivalence of (4.9) to other common formulas in the literature is proved in [5, Thm. 2.1]. crossing form $\mathfrak{q}$ on $\mathcal{L}(t_{0})$ by
$\mathfrak{q}(v,v)=\left<x,\big(X^{*}(t_{0})Y^{\prime}(t_{0})-Y^{*}(t_{0})X^{\prime}(t_{0})\big)x\right>_{\mathbb{C}^{n}},\qquad\text{where }v=\begin{pmatrix}X(t_{0})x\\ Y(t_{0})x\end{pmatrix}$. (4.9)
@@
There are two ways of viewing the formula (4.10), depending on how one defines the Maslov index. One approach is to define the Maslov index for any continuous path via the spectral flow of a certain family of unitary operators. For a regular path one then recovers (4.10) as a special case; see [7, 12] for details. Alternately, one can take (4.10) as the definition of the Maslov index for a regular path, then define the Maslov index of a continuous path by first deforming it to a regular path and then arguing that the resulting index is independent of the choice of deformation, as in [21].
Theorem 4.5.
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\min\left\{\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(t),\mathcal{M}\big)\Big|\ \begin{matrix}\mathcal{L}(0)=\mathcal{L}_{0},\mathcal{L}(1)=\mathcal{L}_{1},\\ t\mapsto\mathcal{L}(t)\text{\ is non-decreasing}\end{matrix}\right\}.$ (4.11)
This theorem highlights the usefulness of the Duistermaat index, since computation of the Maslov index is far more difficult than applying an algebraic identity such as (1.13).
@@
Let $Q\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ be an arbitrary orthogonal projector with $\dim\ker Q=\dim\mathcal{L}_{0}\cap\mathcal{L}_{1}$. Since $\mathcal{G}_{0}\cap\mathcal{G}_{Q}$ and $\mathcal{L}_{0}\cap\mathcal{L}_{1}$ have the same dimension, by Corollary 4.3 there exists a symplectic transformation $\mathsf{S}$ so that $\mathsf{S}\mathcal{L}_{0}=\mathcal{G}_{0}$ and $\mathsf{S}\mathcal{L}_{1}=\mathcal{G}_{Q}$. Since both quantities in (4.12) are symplectic invariants, without loss of generality we can assume that $\mathsf{S}$ is identity. That is, it suffices to find a non-decreasing path $\mathcal{L}(\cdot)$ with $\mathcal{L}(0)=\mathcal{G}_{0}$, $\mathcal{L}(1)=\mathcal{G}_{Q}$ and $\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{Q},\mathcal{M})=\operatorname{Mas}\!\big(\mathcal{L}(t),\mathcal{M}\big)$ for any Lagrangian plane $\mathcal{M}$. We claim that $\mathcal{L}(t)=\mathcal{G}_{tQ}$ has this property.
Representing $\mathcal{M}$ using a Hermitian matrix $M$ on $\operatorname{dom}\mathcal{M}$, as in (3.15), and letting $P$ denote the orthogonal projection onto $\operatorname{dom}\mathcal{M}$, we use the cocycle property (1.10) to obtain
@@
$\mathcal{L}(s,t)\cap\mathcal{M}=\big\{(x,Mx):x\in\operatorname{dom}\mathcal{L}\text{ and }Mx=tPQx-sx\big\}.$ (4.16)
@@
We now evaluate the remaining term, $\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)$. The $t$-crossing form evaluates to $\left<x,Qx\right>_{\mathbb{C}^{n}}$, which we claim to be strictly positive on $\mathcal{L}(s_{*},t)\cap\mathcal{M}$ whenever this intersection is non-trivial. Indeed, the form is obviously non-negative and if $Qx=0$ for some nonzero $x\in\operatorname{dom}\mathcal{L}$, then (4.16) implies $Mx=-s_{*}x$, which is excluded by our condition that $M$ has no spectrum in $s_{*},0)$.
It follows that each $n_{+}$ term in (4.10) is simply the dimension of the space $\mathcal{L}(s_{*},t)\cap\mathcal{M}$,
$\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)=\sum_{t\in[0,1)}\dim\ker\big(M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I\big).$ (4.17)
Since the eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I$ are decreasing in $t$, the sum in (4.17) is equal to the number of eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I$ that cross 0 to become negative, thus
$\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)=n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I\big)-n_{-}(M+s_{*}I).$ (4.18)
@@
$n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I\big)=n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}\big),\qquad n_{-}(M+s_{*}I)=n_{-}(M),$ (4.19)
@@
If $\ker M\cap\ker Q$ is nontrivial, then this subspace is contained in $\mathcal{L}(0,t)\cap\mathcal{M}$ for all $t$. This means the crossings are not isolated and the crossing form is degenerate, so (4.10) does not apply. This is the reason for the introduction of $s$ in the proof above. A similar homotopy argument was used in [21] to prove the “zero axiom” of the Maslov index, which says that a path $\mathcal{L}(t)$ for which $\dim(\mathcal{L}(t)\cap\mathcal{M})$ is constant dimension has $\operatorname{Mas}\!\big(\mathcal{L}(t),\mathcal{M}\big)=0$.
As a corollary of Lemma 4.6 we obtain the Zhou–Wu–Zhu identity [24, Thm. 1.1], which has numerous applications in the analysis of differential equations [2, 16, 11, 5, 18].
@@
For any path $\mathcal{L}\colon[0,1]\to\Lambda(n)$ and any Lagrangian planes $\mathcal{M}_{1}$, $\mathcal{M}_{2}$,
@@
Equation (4.20) now follows by application of (3.2), and equation (4.21) follows from the cocycle property (1.10). ∎
@@
because the path $\mathcal{L}(\cdot)$ is non-decreasing. Lemma 4.6 concludes the proof. ∎
@@ 5. The Morse index of a difference of Hermitian matrices
We start by proving Theorem 1.4, which asserts that
for any $n\times n$ Hermitian matrices $A$, $B$ and $C$.
Proof of Theorem 1.4.
We first use the cocycle property to write
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},\mathcal{G}_{C})=\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},{\mathcal{G}_{\infty}})-\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{C},{\mathcal{G}_{\infty}})+\operatorname{\iota}(\mathcal{G}_{B},\mathcal{G}_{C},{\mathcal{G}_{\infty}}).$ (5.2)
Next, we observe that for any Hermitian $A$ and $B$, the symplectic transformation
$\mathsf{S}:=\begin{pmatrix}I&0\\ A&I\end{pmatrix}$ (5.3)
maps ${\mathcal{G}_{0}}\mapsto\mathcal{G}_{A}$, $\mathcal{G}_{B-A}\mapsto\mathcal{G}_{B}$, and ${\mathcal{G}_{\infty}}\mapsto{\mathcal{G}_{\infty}}$. Symplectic invariance and the normalization condition (1.8) together imply
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},{\mathcal{G}_{\infty}})=\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{B-A},{\mathcal{G}_{\infty}})=n_{-}(B-A).$ (5.4)
Applying this to all three terms on the right-hand side of (5.2) completes the proof. ∎
@@
and observe that for invertible $A$ we have
@@
To establish (5.6) under the assumption that $B$ is invertible, we use Theorem 1.4, equation (5.7), an analogue of (5.8) with AA and $B$ swapped, and, finally, (3.19). ∎
@@
Similar computations lead to a much simpler proof of [5, Prop. 6.1], which we give below.
@@
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=n_{-}\left((\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}-(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}\right).$ (5.12)
@@
We have the following chain of equalities
@@
In the first step we applied (3.5) and used $\mathcal{L}_{1}\cap\mathcal{L}_{3}=\mathcal{L}_{2}\cap\mathcal{L}_{3}=\{0\}$. In the second step we used (3.16), which is allowed because $\mathcal{L}_{3}$ is transversal to $\mathcal{G}_{\infty}$} and therefore is a graph of a matrix. In the third step we used (3.17). We remark that because $\mathcal{L}_{3}$ is transversal to $\mathcal{L}_{1}$, $\mathcal{L}_{1}-\mathcal{L}_{3}$ is transversal to $\mathcal{G}_{0}$} and thus $(\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}$ is transversal to $\mathcal{G}_{\infty}$}. Similarly, $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ is also a graph. Therefore, in the last step, we applied (3.16) to subtract $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ and used the normalization condition (1.8). ∎
@@
which establishes (1.1). Proposition 5.1 can be proved analogously, however, this requires an appropriate extension of the Haynsworth formula such as [19, Thm. 6.1] or [4, Thm. A.1].

### v3

@@ 4.2. The Maslov index
Suppose $\mathcal{L}\colon[0,1]\to\Lambda(n)$ is a differentiable path of Lagrangian planes, given by a frame $\big(X(t),Y(t)\big)$. For $t_{0}\in[0,1]$ and $v\in\mathcal{L}(t_{0})$ we define the [^4] The equivalence of (4.8) to other common formulas in the literature is proved in [15, Thm. 2.1]. crossing form $\mathfrak{q}$ on $\mathcal{L}(t_{0})$ by
$\mathfrak{q}(v,v)=\left<x,\big(X^{*}(t_{0})Y^{\prime}(t_{0})-Y^{*}(t_{0})X^{\prime}(t_{0})\big)x\right>_{\mathbb{C}^{n}},\quad\text{where }v=\big(X(t_{0})x,Y(t_{0})x\big).$ (4.8)
@@
There are two ways of viewing the formula (4.9), depending on how one defines the Maslov index. One approach is to define the Maslov index for any continuous path via the spectral flow of a certain family of unitary operators. For a regular path one then recovers (4.9) as a special case; see [16, 26] for details. Alternately, one can take (4.9) as the definition of the Maslov index for a regular path, then define the Maslov index of a continuous path by first deforming it to a regular path and then arguing that the resulting index is independent of the choice of deformation, as in [48]. $\Diamond$
@@
Let $Q\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ be an arbitrary orthogonal projector with $\dim\ker Q=\dim\mathcal{L}_{0}\cap\mathcal{L}_{1}$. Since $\mathcal{G}_{0}\cap\mathcal{G}_{Q}$ and $\mathcal{L}_{0}\cap\mathcal{L}_{1}$ have the same dimension, there exists a symplectic transformation $\mathsf{S}$ so that $\mathsf{S}\mathcal{L}_{0}=\mathcal{G}_{0}$ and $\mathsf{S}\mathcal{L}_{1}=\mathcal{G}_{Q}$ (see, for example, [46, Exercise 8.1]). Since both quantities in (4.10) are symplectic invariants, without loss of generality we can assume that $\mathsf{S}$ is the identity. That is, it suffices to find a non-decreasing path $\mathcal{L}(\cdot)$ with $\mathcal{L}(0)=\mathcal{G}_{0}$, $\mathcal{L}(1)=\mathcal{G}_{Q}$ and $\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{Q},\mathcal{M})=\operatorname{Mas}\!\big(\mathcal{L}(t),\mathcal{M}\big)$ for any Lagrangian plane $\mathcal{M}$. We claim that $\mathcal{L}(t)=\mathcal{G}_{tQ}$ has this property.
Representing $\mathcal{M}$ using a Hermitian matrix $M$ on $\operatorname{dom}\mathcal{M}$, as in (3.17), and letting $Q_{\operatorname{dom}\mathcal{M}}$ denote the compression of $Q$ to $\operatorname{dom}\mathcal{M}$ (see (3.22)), we use the cocycle property (1.8) to obtain
@@
$\mathcal{L}(s,t)\cap\mathcal{M}=\big\{(x,Mx):x\in\operatorname{dom}\mathcal{M}\text{ and }Mx=tQx-sx\big\}.$ (4.14)
@@
We now evaluate the remaining term, $\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)$. The $t$-crossing form evaluates to $\left<x,Qx\right>_{\mathbb{C}^{n}}$, which we claim to be strictly positive on $\mathcal{L}(s_{*},t)\cap\mathcal{M}$ whenever this intersection is non-trivial. Indeed, the form is obviously non-negative and if $Qx=0$ for some nonzero $x\in\operatorname{dom}\mathcal{M}$, then (4.14) implies $Mx=-s_{*}x$, which is excluded by our condition that $M$ has no spectrum in $s_{*},0)$, thus proving the claim. The same argument shows that any eigenvalue curve $\lambda(t)$ for the matrix $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I_{\operatorname{dom}\mathcal{M}}$ has $\lambda^{\prime}(t)\leq 0$ in general and $\lambda^{\prime}(t_{0})<0$ at any time $t_{0}$ for which $\lambda(t_{0})=0$.
It follows that each $n_{+}$ term in (4.9) equals the dimension of the space $\mathcal{L}(s_{*},t)\cap\mathcal{M}$, so
$\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)=\sum_{t\in[0,1)}\dim\ker\big(M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I_{\operatorname{dom}\mathcal{M}}\big).$ (4.15)
Since the eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I_{\operatorname{dom}\mathcal{M}}$ are strictly decreasing in $t$ when they equal zero, the sum in (4.15) is equal to the number of eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I_{\operatorname{dom}\mathcal{M}}$ that pass through zero (and hence become negative) at some $t\in[0,1)$, thus
$\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(s_{*},t),\mathcal{M}\big)=n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I\big)-n_{-}(M+s_{*}I_{\operatorname{dom}\mathcal{M}}).$ (4.16)
@@
$n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I_{\operatorname{dom}\mathcal{M}}\big)=n_{-}\big(M-Q_{\operatorname{dom}\mathcal{M}}\big),\qquad n_{-}(M+s_{*}I)=n_{-}(M),$ (4.17)
@@
We note that $\ker M\cap\ker Q\not=\{0\}$ implies $\mathcal{L}(0,t)\cap\mathcal{M}\not=\{0\}$ for all $t$, because $(\ker M\cap\ker Q)\oplus\{0\}\subset\mathcal{L}(0,t)\cap\mathcal{M}$. This means the crossings are not isolated and the crossing form is degenerate, so (4.9) does not apply. This is the reason for the introduction of $s$ in the proof above. A similar homotopy argument was used in [48] to prove the “zero axiom” of the Maslov index, which says that a path $\mathcal{L}(t)$ for which $\dim(\mathcal{L}(t)\cap\mathcal{M})$ is constant has $\operatorname{Mas}\!\big(\mathcal{L}(t),\mathcal{M}\big)=0$. $\Diamond$
As a corollary of Lemma 4.4 we obtain the Zhou–Wu–Zhu identity [51, Thm. 1.1], which has numerous applications in the analysis of differential equations [8, 32, 25, 15, 44].
@@
For any path $\mathcal{L}\colon[0,1]\to\Lambda(n)$ and any Lagrangian planes $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$,
@@
Equation (4.18) now follows from (3.2), and equation (4.19) follows from the cocycle property (1.8). ∎
We are now ready to proved the relationship (1.14) between the Duistermaat and Maslov indices that was claimed in the introduction.
@@
because the path $\mathcal{L}(\cdot)$ is non-decreasing. Lemma 4.4 shows that the upper bound is achievable. Finally, once the path $\mathcal{L}(t)$ achieves equality for some $\mathcal{M}$, i.e.,
$\operatorname{\iota}\!\big(\mathcal{L}(0),\mathcal{L}(1),\mathcal{M}\big)=\operatorname{Mas}_{[0,1]}\!\big(\mathcal{L}(t),\mathcal{M}\big),$ (4.23)
the same equality holds for any other $\mathcal{M}$ by (4.18). ∎
@@ 5. The Morse index of a difference of Hermitian matrices
and observe that for invertible AA we have
@@
To establish (5.2) under the assumption that $B$ is invertible, we use Theorem 1.4, equation (5.3), an analogue of (5.4) with AA and $B$ swapped and, finally, (3.21). ∎
@@
Similar computations lead to a much simpler proof of [15, Prop. 6.1], which we give below.
@@
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=n_{-}\left((\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}-(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}\right);$ (5.8)
the relation in the right-hand side is the graph of a matrix and $n_{-}$ denotes its Morse index.
@@
We have the following chain of equalities:
@@
In the first step we applied (3.5) and used $\mathcal{L}_{1}\cap\mathcal{L}_{3}=\mathcal{L}_{2}\cap\mathcal{L}_{3}=\{0\}$. In the second step we used (3.18), which is allowed because $\mathcal{L}_{3}$ is transversal to $\mathcal{G}_{\infty}$} and therefore is the graph of a matrix. In the third step we used (3.19). We remark that because $\mathcal{L}_{3}$ is transversal to $\mathcal{L}_{1}$, $\mathcal{L}_{1}-\mathcal{L}_{3}$ is transversal to $\mathcal{G}_{0}$} and thus $(\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}$ is transversal to $\mathcal{G}_{\infty}$}. Similarly, $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ is also a graph. Therefore, in the last step, we used (3.18) to subtract $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ and used the normalization condition (1.6). ∎
@@
which establishes (1.11). Proposition 5.1 can be proved analogously; this requires a suitable extension of the Haynsworth formula such as [45, Thm. 6.1] or [14, Thm. A.1]. $\Diamond$